A ball is thrown vertically upward from the foot of a tower. it crosses the top of the tower twice after an interval of 4s and reaches the foot of the tower 8s after its was thrown. what is the height of the tower

Grade:12th

1 Answers

Arun
6 years ago
Dear Samresh
 
time taken to go upto U and then come back to the ground required 8 sec.
So 4 s up then 4 s in coming down.
the time taken between climbing from T to U and then coming to T again(crossing T twice) =4 s
so time taken to go upto T from the ground=(8 - 4)/2=2s
now if we consider the journey from ground to U:
0=u-4x10 .......  putting in 1st eq.
u=40 m/s
now tower height:H=
H=40x2-1/2  x10x2x2.....putting in second eq
H=60 m
 
Regards
Arun (askIITians forum expert)